OpenStudy (anonymous):

what is the measure of the slope that passes through (2,6) and (5,-3)? A.-3 B.-1/3 C.1/3 D.3

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

numerator = y1 -y2 = 6- (-3) = ... ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

after this one can you help me with 4 more?

hartnn (hartnn):

no, its not 3

OpenStudy (anonymous):

-3?

hartnn (hartnn):

\(6-(-3) = 6+3 = ...\)

OpenStudy (anonymous):

oh 9

hartnn (hartnn):

yes denominator = x1 -x2 =2-5 = ...?

OpenStudy (anonymous):

-3

hartnn (hartnn):

yes slope = m = 9/(-3) =..

OpenStudy (anonymous):

so it would be d 3

hartnn (hartnn):

no... m = 9/(-3) = -9/3 = -3

OpenStudy (anonymous):

ok

hartnn (hartnn):

did you understand the procedure ?

hartnn (hartnn):

first find difference in y co-ordinates numerator = y1 -y2 then find difference in x co-ordinates denominator = x1 -x2 slope m = numerator/ denominator = (y1-y2)/(x1-x2)

OpenStudy (anonymous):

What is the slope of the line? that passes through (4,1) and (-7,-2) A.1 1/3 B.3/11 C.-3/11 D.-11/3

hartnn (hartnn):

try by yourself first

OpenStudy (anonymous):

i think it would be maybe b?

hartnn (hartnn):

B is correct :)

OpenStudy (anonymous):

What is the slope of the line that passes through the points (2, 8) and (6, 12)? A. –5 B. –1 C. 1 D. 5

OpenStudy (anonymous):

i think it would be b?

hartnn (hartnn):

nopes

hartnn (hartnn):

numerator = y1-y2 = ... ?

OpenStudy (anonymous):

c

hartnn (hartnn):

yes, C

OpenStudy (anonymous):

What is the slope of the line that passes through the points (16, –1) and (–4, 10)? A.-20/11 B.-11/20 C.11/20 D.20/11

OpenStudy (anonymous):

b?

hartnn (hartnn):

correct!

OpenStudy (anonymous):

What is the slope of the line that passes through the points (–2, 15) and (–8, –5)? A.-10/3 B.-5/2 C.5/2 D.10/3

OpenStudy (anonymous):

i think it is d?

hartnn (hartnn):

good! D is correct :)

OpenStudy (anonymous):

can you help me with 9 other questions?

OpenStudy (anonymous):

What is the slope of the graph of 4x – 16y = 15? A.-4 B.-1/4 C.1/4 D.4

hartnn (hartnn):

ask in a new post, i am unavailable, others will help you :)